Given sin theta = sqrt(23)/12 and tan theta = sqrt(23)/11 find cos theta
Answer by Alan3354(69443)

Given sin theta = sqrt(23)/12 and tan theta = sqrt(23)/11 find cos theta
===============
tan = sin/cos
cos = sin/tan
cos = (sqrt(23)/12)/(sqrt(23)/11)
= 11/12